### v2

@@ Abstract
Inferential models (IMs) offer provably reliable, data-driven, possibilistic statistical inference. But despite IMs’ theoretical and foundational advantages, efficient computation is often a challenge. This paper presents a simple and powerful numerical strategy for approximating the IM’s possibility contour, or at least its $\alpha$-cut for a specified $\alpha\in(0,1)$. Our proposal starts with the specification a parametric family that, in a certain sense, approximately covers the credal set associated with the IM’s possibility measure. Then the parameters of that parametric family are tuned in such a way that the family’s $100(1-\alpha)\%$ credible set roughly matches the IM contour’s $\alpha$-cut. This is reminiscent of the variational approximations now widely used in Bayesian statistics, hence the name variational-like IM approximation.
@@ 1 Introduction
While Bayesian and frequentist are currently the two mainstream schools of thought in statistical inference, these are not the only perspectives. For example, Dempster–Shafer theory originated as an improvement to and generalization of both Bayesian inference and Fisher’s fiducial argument. Of particular interest to us here are the recent advances in inferential models (IMs, Martin and Liu, 2013; Martin and Liu, 2015b ; Martin, 2021a ; Martin, 2022b ), a framework that offers Bayesian-like, data-dependent, possibilistic quantification of uncertainty about unknowns but with built-in, frequentist-like reliability guarantees. IMs and other new/non-traditional frameworks are currently facing the same computational challenges that Bayesians faced years ago. That is, we know what we want to compute and why, but we are currently lacking the tools to do so efficiently. While traditional Monte Carlo methods are still useful (see Section 2), the imprecision that is key to the IM’s reliability guarantees implies that Monte Carlo methods alone are not enough. In contrast to Lindley’s prediction, for Efron’s speculation about fiducial and related methods—“Maybe Fisher’s biggest blunder will become a big hit in the 21st century!” (Efron, 1998)—to come true, imprecision-accommodating advances in Monte Carlo computations are imperative. The present paper’s contribution is in this direction.
@@
The remainder of the paper is organized as follows. After some background on possibilsitic IMs and their properties in Section 2, we present our basic but very general variational-like IM approximation in Section 3, which relies on a combination of Monte Carlo and stochastic approximation to tune the index of the posited parametric family of approximations. This is ideally suited for statistical inference problems involving low-dimensional parameters, and we offer a few such illustrations in Section 4. A more sophisticated version of the proposed variational-like approximation is presented in Section 5, which is more appropriate when the problem at hand is of higher dimension, but it is suited primarily for Gaussian approximation families—which is no practical restriction. With a more structured approximation, we can lower the computational cost by reducing the number of Monte Carlo evaluations of the IM contour. This is illustrated in several examples, including a relatively high-dimensional problem with lasso penalization and in problems involving nuisance parameters in parametric, nonparametric, and semiparametric models, respectively. The paper concludes in Section 7 with a concise summary and a discussion of some practically relevant extensions.
@@ 2 Background on possibilistic IMs
Consider a parametric statistical model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ indexed by a parameter space $\mathbb{T}\subseteq\mathbb{R}^{d}$. Examples include $\mathsf{P}_{\theta}={\sf Bin}(n,\theta)$, $\mathsf{P}_{\theta}={\sf N}(\theta,1)$, $\mathsf{P}_{\theta}={\sf Gamma}(\alpha,\beta)$ with $\theta=(\alpha,\beta)$, and many others; see Section 4. Nonparametric problems—see Cella, (2024), Cella and Martin, (2022), and Martin, 2023b (, Sec. 5)—can also be handled, but we postpone discussion of this until Section 6. Suppose that the observable data $X^{n}=(X_{1},\ldots,X_{n})$ consists of iid samples from the distribution $\mathsf{P}_{\Theta}$, where $\Theta\in\mathbb{T}$ is the unknown/uncertain “true value.” The model and observed data $X^{n}=x^{n}$ together determine a likelihood function $\theta\mapsto L_{x^{n}}(\theta)$ and a corresponding relative likelihood function
@@
Throughout we will implicitly assume that the denominator is finite, i.e., that the likelihood function is bounded for almost all $x^{n}$.
@@
A corresponding necessity measure, or lower probability, is defined via conjugacy as $\underline{\Pi}_{x^{n}}(H)=1-\overline{\Pi}_{x^{n}}(H^{c})$. This measure plays a crucial role in quantifying the support for hypotheses of interest (Cella and Martin, 2023); see Example 5.
An essential feature of this IM construction is its so-called validity property:
@@
This has a number of important consequences. First, (2) is exactly the calibration property required to ensure a “confidence” connection that the relative likelihood on its own misses. In particular, (2) immediately implies that the set
@@
The above-described operations applied to the IM’s contour function (1), to obtain upper probabilities or to eliminate nuisance parameters, are all special cases of Choquet integrals (e.g. Troffaes and de Cooman, 2014, App. C). These more general Choquet integrals can be statistically relevant in formal decision-making contexts (Martin, 2021b ), etc. For the sake of space, we will not dig into these details here, but the fact that there are non-trivial operations to be carried out even after the IM’s contour function has been obtained provides genuine, practical motivation for us to find the simplest and most efficient contour approximation as possible.
While the IM construction and properties are relatively simple, computation can be a challenge. The key observation is that we rarely have the sampling distribution of the relative likelihood $R(X^{n},\theta)$, under $\mathsf{P}_{\theta}$, available in closed form to facilitate exact computation of $\pi_{x^{n}}(\theta)$ for any $\theta$. So, instead, the go-to strategy is to approximate that sampling distribution using Monte Carlo at each value of $\theta$ on a sufficiently fine grid (e.g., Martin, 2022b ; Hose et al., 2022). That is, the possibility contour is approximated as
$\pi_{x^{n}}(\theta)\approx\frac{1}{M}\sum_{m=1}^{M}1\{R(X_{m,\theta}^{n},\theta)\leq R(x^{n},\theta)\},\quad\theta\in\mathbb{T},$
where $1(\cdot)$ is the indicator function and $X_{m,\theta}^{n}$ consists of $n$ iid samples from $\mathsf{P}_{\theta}$ for $m=1,\ldots,M$. The above computation is feasible at one or a few different $\theta$ values, but it is common that this needs to be carried out over a sufficiently fine grid covering the (relevant area) of the often multi-dimensional parameter space $\mathbb{T}$. For example, the confidence set in (3) requires that we can solve the equation $\pi_{x^{n}}(\theta)=\alpha$, or at least find which $\theta$’s satisfy $\pi_{x^{n}}(\theta)\geq\alpha$, and a naive approach is to compute the contour over a huge grid and then keep those that (approximately) solve the aforementioned equation. This amounts to lots of wasted computations. More generally, the relevant summaries of the IM output involve optimization of the contour, and carrying out this optimization numerically requires many contour function evaluations. Simple tweaks to this most-naive approach can be employed in certain cases, e.g., importance sampling, but these adjustments require problem-specific considerations and they cannot be expected to offer substantial improvements in computational efficiency. This is a serious bottleneck, so new and not-so-naive computational strategies are desperately needed.
@@ 3 Basic variational-like IMs
The Monte Carlo-based strategy reviewed above is not the only way to numerically approximate the possibilistic IM. Another option is an analytical “Gaussian” approximation (see below) based on the available large-sample theory (Martin and Williams, 2024). The goal here is to strike a balance between the (more-or-less) exact but expensive Monte Carlo-based approximation and the rough but cheap large-sample approximation. To strike this balance, we choose to focus on a specific feature of the possibilistic IM’s output, namely, the confidence sets $C_{\alpha}(x^{n})$ in (3), and choose the approximation such that it at least matches the given confidence set exactly. Our specific proposal resembles the variational approximations that are now widely used in Bayesian analysis, where first a relatively simple family of candidate probability distributions is specified and then the approximation is obtained by finding the candidate that minimizes (an upper bound on) the distance/divergence from the exact posterior distribution. Our approach differs in the sense that we are aiming to approximate a possibility measure by (the probability-to-possibility transform applied to) an appropriately chosen probability distribution.
@@
(Where convenient, we will replace the subscript “$x^{n}$” by a subscript “$n$”—e.g., $\mathsf{Q}_{n}$ and $\overline{\Pi}_{n}$ instead of $\mathsf{Q}_{x^{n}}$ and $\overline{\Pi}_{x^{n}}$—to simplify the notation in what follows.) That is, a data-dependent probability $\mathsf{Q}_{n}$ is consistent with $\overline{\Pi}_{n}$ if and only if, for each $\alpha\in[0,1]$, it assigns at least $1-\alpha$ mass to the IM’s confidence set $C_{\alpha}(x^{n})$ in (3). Furthermore, the “best” inner probabilistic approximation of the possibilistic IM, if it exists, corresponds to a $\mathsf{Q}_{n}^{\star}$ such that $\mathsf{Q}_{n}^{\star}\{C_{\alpha}(x^{n})\}=1-\alpha$ for all $\alpha\in[0,1]$. For a certain special class of statistical models, Martin, 2023a showed that this best inner approximation corresponds to Fisher’s fiducial distribution and the default-prior Bayesian posterior distribution. Beyond this special class of models, however, it is not clear if a best inner approximation exists and, if so, how to find it. A less ambitious goal is to find, for a fixed choice of $\alpha$, a probability distribution $\mathsf{Q}_{n,\alpha}^{\star}$ such that
@@
$\hat{f}_{\alpha}(\xi)=\frac{1}{M}\sum_{m=1}^{M}1\{\pi_{x^{n}}(\Theta_{m}^{\xi})>\alpha\}-(1-\alpha),$ (8)
@@
It is well-known that, under certain mild conditions, the sequence $(\xi^{(t)}:t\geq 0)$ as defined above converges in probability to the root of $f_{\alpha}$ in (7). If $\hat{\xi}$ is the value returned when the algorithm reaches practical convergence, e.g., when $|\hat{f}_{\alpha}(\xi^{(t)})|$ or the change $|\xi^{(t+1)}-\xi^{(t)}|$ is smaller than some specified threshold, then we set $\widehat{\mathsf{Q}}_{n,\alpha}=\mathsf{Q}_{n,\alpha}^{\hat{\xi}}$. This distribution should be at least a decent approximation of the IM possibility measure’s inner approximation, i.e., the “most diffuse” member of $\mathscr{C}(\overline{\Pi}_{x^{n}})$. Consequently, the probability-to-possibility transform in (1) applied to $\widehat{\mathsf{Q}}_{x^{n},\alpha}$ should be a decent approximation of the exact possibilistic IM contour $\pi_{x^{n}}$, at least in terms of their upper $\alpha$-cuts. The illustrations presented in Section 4 below show that our proposed approximation is much better than “decent” in cases where we can visually compare with the exact IM contour.
@@ 4 Numerical illustrations
Suppose the data is $X\sim{\sf Bin}(n,\Theta)$ with probability mass function $p_{\theta}$. The exact IM possibility contour based on $X=x$ is
$\pi_{x}(\theta)=\sum_{s=0}^{n}1\{R(s,\theta)\leq R(x,\theta)\}\,p_{\theta}(s),$ (10)
where $R(s,\theta)$ is the binomial relative likelihood of $\theta$ for $s$ successes:
$R(s,\theta)=\Bigl{(}\frac{n\theta}{s}\Bigr{)}^{s}\Bigl{(}\frac{n-n\theta}{n-s}\Bigr{)}^{n-s},\quad\theta\in[0,1].$
The right-hand side of (10) can be evaluated numerically without Monte Carlo. This and the contour corresponding to the proposed Gaussian-based variational approximation are shown in Figure 1(a) for $n=15$ and $x=6$ successes. Note that the two contours closely agree, especially at the level $\alpha=0.1$ specifically targeted.
@@
Inference on the unknown correlation $\Theta$ is a surprisingly challenging problem (e.g., Basu, 1964; Reid, 2003; Martin and Liu, 2015a ). Indeed, despite being a scalar exponential family model, it does not have a one-dimensional sufficient statistic and, moreover, there are a variety of different ancillary statistics on which to condition, leading to different solutions. From a computational perspective, it is worth noting that the maximum likelihood estimation is not available in closed-form—in particular, it is not the sample correlation coefficient. That this optimization is non-trivial adds to the burden of a naive Monte Carlo IM construction. Figure 1(b) shows the exact IM contour, based on a naive Monte Carlo implementation of (1), which is rather expensive, and the approximation for simulated data of size $n=50$ with true correlation 0.5. The exact contour has some asymmetry that the normal approximation cannot perfectly accommodate, but it makes up for this imperfection with a slightly wider upper 0.1-level set.
@@
where $F(z)=(1+e^{-z})^{-1}$ is the logistic distribution function. The corresponding likelihood cannot be maximized in closed-form, but doing so numerically is routine. The maximum likelihood estimator and the corresponding observed information matrix lead to the asymptotically valid inference reported by standard statistical software packages. For exact inference, however, the computational burden is heavier: evaluating the validified relative likelihood over a sufficiently fine grid of $\theta$ values is rather expensive. Figure 1(c) presents the $0.1$-level set of the exact IM possibility contour for the regression coefficients, based on a naive Monte Carlo implementation of (1), alongside the proposed variational approximation. The latter’s computational cost is a small fraction of the former’s, and yet the two contours line up almost perfectly.
@@
Poisson log-linear models are widely used to analyze the relationship between a count-based discrete response variable and a set of fixed explanatory variables; even if the explanatory variables are not fixed by design, it is rare that their distribution would depend on any of the relevant parameters, so they are ancillary statistics and it is customary to condition on their observed values. Let $X_{i}$ represent the response variable, and $z_{i1},\ldots,z_{ip}$ denote the $p$ explanatory variables for the $i^{th}$ observation in a sample, $i=1,\ldots,n$. The Poisson log-linear model assumes that $X_{i}\sim{\sf Pois}(\Lambda_{i})$, where
$\log\Lambda_{i}=\Theta_{0}+\Theta_{1}z_{i1}+\Theta_{2}z_{i2}+\ldots+\Theta_{p}z_{ip}.$
The goal is to quantify uncertainty about the regression coefficients $\Theta_{0},\Theta_{1},\ldots,\Theta_{p}$. Here, we will achieve this through the IMs framework.

### v3

@@ Abstract
Inferential models (IMs) offer provably reliable, data-driven, possibilistic statistical inference. But despite the IM framework’s theoretical and foundational advantages, efficient computation is a challenge. This paper presents a simple yet powerful numerical strategy for approximating the IM’s possibility contour, or at least its $\alpha$-cut for a specified $\alpha\in(0,1)$. Our proposal starts with the specification of a parametric family that, in a certain sense, approximately covers the credal set associated with the IM’s possibility measure. Akin to variational inference, we then propose to tune the parameters of that parametric family so that its $100(1-\alpha)\%$ credible set roughly matches the IM contour’s $\alpha$-cut. This parametric $\alpha$-cut matching strategy implies a full approximation to the IM’s possibility contour at a fraction of the computational cost associated with previous strategies.
@@ 1 Introduction
While Bayesian and frequentist are currently the two mainstream schools of thought in statistical inference, these are not the only perspectives. For example, Dempster–Shafer theory originated as an improvement to and generalization of both Bayesian inference and Fisher’s fiducial argument. Of particular interest to us here are the recent advances in inferential models (IMs, Martin and Liu, 2013; Martin and Liu, 2015b ; Martin, 2021a ; Martin, 2022b ), a framework that offers Bayesian-like, data-dependent, possibilistic quantification of uncertainty about unknowns but with built-in, frequentist-like reliability guarantees. IMs and other new/non-traditional frameworks are currently facing the same computational challenges that Bayesians faced years ago. That is, we know what we want to compute and why, but we are currently lacking the tools to do so efficiently. While traditional Monte Carlo methods are still useful (see Section 2), the imprecision that is key to the IM’s reliability guarantees implies that Monte Carlo methods alone are not enough. In contrast to Lindley’s, for Efron’s prediction about fiducial-like methods—“Maybe Fisher’s biggest blunder will become a big hit in the 21st century!” (Efron, 1998)—to come true, imprecision-accommodating advances in Monte Carlo computations are imperative. The present paper’s contribution is in this general direction, offering a new IM approximation and an efficient method for computing that approximation.
@@
The remainder of the paper is organized as follows. After some background on possibilistic IMs and their properties in Section 2, we present our basic but very general variational-like IM approximation in Section 3, which relies on a combination of Monte Carlo and stochastic approximation to tune the index of the posited parametric family of approximations. This is ideally suited for statistical inference problems involving low-dimensional parameters, and we offer a few such illustrations. A more sophisticated version of the proposed variational-like approximation is presented in Section 4, which is more appropriate when the problem at hand is of higher dimension, but it is suited primarily for Gaussian approximation families—which is no practical restriction. With a more structured approximation, we can lower the computational cost by reducing the number of Monte Carlo evaluations of the IM contour. This is illustrated in several examples, including a relatively high-dimensional problem with lasso penalization and in problems involving nuisance parameters in parametric, nonparametric, and semiparametric models, respectively. We emphasize that this paper is not about a new IM construction. Our contribution is a new and computationally efficient approximation to the IM solution that has been developed and studied elsewhere in the literature. So, rather than comparing our IM approximation against other methods (e.g., Bayesian), the illustrations presented here focus on the quality of our newly proposed approximation and showing that it closely matches the often substantially more expensive exact IM solution that has already been studied. If the proposed approximation is accurate, then the comparisons between the IM solution and other methods—see, e.g., Cahoon and Martin, (2021, 2020); Martin, (2015); Martin, 2022b ; Martin, 2023b ; Cella, (2024); Cella and Martin, 2022a ; Cella and Martin, 2022b —would be the same whether using the old and expensive computational strategy or the new and more efficient one proposed here. We do, however, include a brief comparison between IMs and the Bayesian approach; see Example 5. The paper concludes in Section 6 with a concise summary and a discussion of some practically relevant extensions.
@@ 2 Background on possibilistic IMs
Consider a parametric statistical model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ indexed by a parameter space $\mathbb{T}\subseteq\mathbb{R}^{d}$. Examples include $\mathsf{P}_{\theta}={\sf Ber}(\theta)$, $\mathsf{P}_{\theta}={\sf N}(\theta,1)$, $\mathsf{P}_{\theta}={\sf Gamma}(\alpha,\beta)$ with $\theta=(\alpha,\beta)$, and many others; see Section 3.3. Nonparametric problems—see Cella, (2024), Cella and Martin, 2022a , and Martin, 2023b (, Sec. 5)—can also be handled, but we postpone discussion of this until Section 5. Suppose that the observable data $X^{n}=(X_{1},\ldots,X_{n})$ consists of iid samples from the distribution $\mathsf{P}_{\Theta}$, where $\Theta\in\mathbb{T}$ is the unknown/uncertain “true value.” The model and observed data $X^{n}=x^{n}$ together determine a likelihood function $\theta\mapsto L_{x^{n}}(\theta)$ and a corresponding relative likelihood function
@@
Throughout we will implicitly assume that the denominator is finite. As a running example, consider a binomial problem where $X^{n}$ represent the outcomes of $n$ independent Bernoulli trials with unknown probability of success $\Theta$. Then the relative likelihood is
$R(x^{n},\theta)=\left(\frac{n\theta}{\sum_{i=1}^{n}x_{i}}\right)^{\sum_{i=1}^{n}x_{i}}\left(\frac{n-n\theta}{n-\sum_{i=1}^{n}x_{i}}\right)^{n-\sum_{i=1}^{n}x_{i}},\quad\theta\in[0,1].$ (1)
@@
A corresponding necessity measure, or lower probability, is defined via conjugacy as $\underline{\Pi}_{x^{n}}(H)=1-\overline{\Pi}_{x^{n}}(H^{c})$. This measure plays a crucial role in quantifying the support for hypotheses of interest (Cella and Martin, 2023); see Example 5. For our binomial running example with relative likelihood (1), the possibility contour based on $X^{n}=x^{n}$ is
$\pi_{x^{n}}(\theta)=\sum_{s=0}^{n}1\{R(s,\theta)\leq R(x^{n},\theta)\}\,p_{\theta}(s),$ (3)
where $1(\cdot)$ is the indicator function and $p_{\theta}$ is the probability mass function of ${\sf Bin}(n,\theta)$. This can be easily evaluated, without Monte Carlo, and it is displayed in Figure 1(a).
An essential feature of above IM construction is its so-called validity property:
@@
The proof of (4) is a simple application of the now-standard probability integral transform (originally due to Fisher, 1973, Sec. 21.1), i.e., if a random variable $Z$ has distribution function $G$, then the random variable $G(Z)$ is stochastically no larger than ${\sf Unif}(0,1)$. The property (4) has a number of important consequences. First, (4) is exactly the calibration property required to ensure a “confidence” connection that the relative likelihood on its own misses. In particular, (4) immediately implies that the set
@@
The above-described operations applied to the IM’s contour function (2), to obtain upper probabilities or to eliminate nuisance parameters, are all special cases of Choquet integrals (e.g. Troffaes and de Cooman, 2014, App. C). These more general Choquet integrals can be statistically relevant in formal decision-making contexts (Martin, 2021b ), etc. That is, if $\ell_{a}(\Theta)$ represents the loss associated with taking action $a$ when the state of the world is $\Theta$, then the possibilistic IM provides an assessment of the risk associated with action $a$ that accounts for uncertainty in $\Theta$, given $x^{n}$, via the Choquet integral:
[…]
Then, for example, one might choose the action $\hat{a}(x^{n})$ that minimizes this upper expected loss. The point is that there are non-trivial operations to be carried out even after the IM’s contour function has been obtained provides genuine, practical motivation for us to find the simplest and most efficient contour approximation as possible.
While the IM construction is conceptually simple and its properties are strong, computation can be a challenge. The issue is that we rarely have the sampling distribution of the relative likelihood $R(X^{n},\theta)$, under $\mathsf{P}_{\theta}$, available in closed form to facilitate exact computation of $\pi_{x^{n}}$. So, instead, the go-to strategy is to approximate that sampling distribution using Monte Carlo at each value of $\theta$ on a sufficiently fine grid (e.g., Martin, 2022b ; Hose et al., 2022). That is, the possibility contour is approximated as
$\pi_{x^{n}}(\theta)\approx\frac{1}{M}\sum_{m=1}^{M}1\{R(X_{m,\theta}^{n},\theta)\leq R(x^{n},\theta)\},\quad\theta\in\mathbb{T},$ (7)
where $X_{m,\theta}^{n}$ consists of $n$ iid samples from $\mathsf{P}_{\theta}$ for $m=1,\ldots,M$. The above computation is feasible at one or a few different $\theta$ values, but it is common that this needs to be carried out over a sufficiently fine grid covering the (relevant area) of the often multi-dimensional parameter space $\mathbb{T}$. For example, the confidence set in (5) requires that we can solve the equation $\pi_{x^{n}}(\theta)=\alpha$, or at least find which $\theta$’s satisfy $\pi_{x^{n}}(\theta)\geq\alpha$, and a naive approach is to compute the contour over a huge grid and then keep those that (approximately) solve the aforementioned equation. The computational complexity associated with this is $O(Mg^{d})$, where $M$ is the Monte Carlo sample size, $d$ is the dimension of the parameter space, and $g$ is number of grid points in each dimension of $\mathbb{T}$; this amounts to lots of wasted computations. More generally, the relevant summaries of the IM output involve optimization of the contour, and carrying out this optimization numerically requires many contour function evaluations. Simple tweaks to this most-naive approach can be employed in certain cases, e.g., importance sampling, but these adjustments require problem-specific considerations and they cannot be expected to offer substantial improvements in computational efficiency. This is a serious bottleneck, so new and not-so-naive computational strategies are desperately needed.
@@ 3 Basic variational-like IMs
3.1 Setup
The Monte Carlo-based strategy reviewed above is not the only way to numerically approximate the possibilistic IM. Another option is an analytical “Gaussian” approximation (see below) based on the available large-sample theory (Martin and Williams, 2025). The goal here is to strike a balance between the (more-or-less) exact but expensive Monte Carlo-based approximation and the rough but cheap large-sample approximation. To strike this balance, we choose to focus on a specific feature of the possibilistic IM’s output, namely, the confidence sets $C_{\alpha}(x^{n})$ in (5), and choose the approximation such that it at least matches the given confidence set exactly. Our specific proposal resembles the variational approximations that are now widely used in Bayesian analysis, where first a relatively simple family of candidate probability distributions is specified and then the approximation is obtained by finding the candidate that minimizes (an upper bound on) the distance/divergence from the exact posterior distribution. Our approach differs in the sense that we are aiming to approximate a possibility measure by (the probability-to-possibility transform applied to) an appropriately chosen probability distribution.
@@ 3.1 Setup
(Where convenient, we will replace the subscript “$x^{n}$” by a subscript “$n$”—e.g., $\mathsf{Q}_{n}$ and $\overline{\Pi}_{n}$ instead of $\mathsf{Q}_{x^{n}}$ and $\overline{\Pi}_{x^{n}}$—to simplify the notation in what follows.) That is, a data-dependent probability $\mathsf{Q}_{n}$ is consistent with $\overline{\Pi}_{n}$ if and only if, for each $\alpha\in[0,1]$, it assigns at least $1-\alpha$ mass to the IM’s confidence set $C_{\alpha}(x^{n})$ in (5). Furthermore, the “best” inner probabilistic approximation of the possibilistic IM, if it exists, corresponds to a $\mathsf{Q}_{n}^{\star}$ such that $\mathsf{Q}_{n}^{\star}\{C_{\alpha}(x^{n})\}=1-\alpha$ for all $\alpha\in[0,1]$. For a certain special class of statistical models, Martin, 2023a showed that this best inner approximation corresponds to Fisher’s fiducial distribution and the default-prior Bayesian posterior distribution. Beyond this special class of models, it is not clear how to find a best inner approximation. A less ambitious goal is to find, for a fixed $\alpha$, a probability distribution $\mathsf{Q}_{n,\alpha}^{\star}$ such that
@@ 3.2 Proposed approximation
$\hat{f}_{\alpha}(\xi)=\frac{1}{K}\sum_{k=1}^{K}1\{\pi_{x^{n}}(\Theta_{k}^{\xi})>\alpha\}-(1-\alpha),$ (11)
@@
Only having an unbiased estimator of the objective function requires some adjustments to the numerical routine. In particular, rather than a Newton–Raphson routine that assumes the function values are noiseless, we must use a stochastic approximation algorithm (e.g., Syring and Martin, 2021, 2019; Martin and Ghosh, 2008; Kushner and Yin, 2003; Robbins and Monro, 1951) that is adapted to noisy function evaluations of $\hat{f}_{\alpha}$. The basic Robbins–Monro algorithm seeks the root of (10) through the updates
@@
Under certain mild conditions, the sequence $(\xi^{(t)}:t\geq 0)$ as defined above converges in probability to the root of $f_{\alpha}$ in (10). If $\hat{\xi}$ is the value returned when the algorithm reaches practical convergence, e.g., when $|\hat{f}_{\alpha}(\xi^{(t)})|$ or the change $|\xi^{(t+1)}-\xi^{(t)}|$ is smaller than some specified threshold, then we set $\widehat{\mathsf{Q}}_{n,\alpha}=\mathsf{Q}_{n,\alpha}^{\hat{\xi}}$. This distribution should be a reasonably accurate approximation of the IM possibility measure’s inner approximation, i.e., the “most diffuse” member of $\mathscr{C}(\overline{\Pi}_{x^{n}})$. Consequently, the probability-to-possibility transform in (2) applied to $\widehat{\mathsf{Q}}_{x^{n},\alpha}$ should be a reasonably accurate approximation of the exact possibilistic IM contour $\pi_{x^{n}}$, at least in terms of their upper $\alpha$-cuts. The illustrations in the following sections confirm that this and the more sophisticated approximation presented in Section 4 below are indeed reasonably accurate.
@@ 3.3 Numerical illustrations
Recall the binomial example explored in Section 2, where $X^{n}$ consists of $n$ iid ${\sf Ber}(\Theta)$ random variables. An expression for the exact IM possibility contour based on $X^{n}=x^{n}$ is given in (3) and this is displayed in Figure 1(a) for $n=15$ and $x^{n}$ such that $\sum_{i=1}^{n}x_{i}=6$. Figure 1(a) also shows the contour corresponding to the proposed Gaussian-based variational approximation. Note that the two contours closely agree, especially at the level $\alpha=0.1$ specifically targeted.
@@
Inference on the unknown correlation $\Theta$ is a surprisingly challenging problem (e.g., Basu, 1964; Reid, 2003; Martin and Liu, 2015a ). Indeed, despite being a scalar exponential family model, it does not have a one-dimensional sufficient statistic and, moreover, there are a variety of different ancillary statistics on which to condition, leading to different solutions. Figure 1(b) shows the exact IM contour, based on a naive Monte Carlo implementation of (2), and the approximation for simulated data of size $n=50$ with true correlation 0.5. The exact contour has some asymmetry that the normal approximation cannot perfectly accommodate, but it makes up for this imperfection with a slightly wider upper 0.1-level set. A quantitative assessment of the computational efficiency and accuracy of the proposed approximation is presented in Example 6 below.
@@
where $F(z)=(1+e^{-z})^{-1}$ is the logistic distribution function. The corresponding likelihood cannot be maximized in closed-form, but doing so numerically is routine. The maximum likelihood estimator and the corresponding observed information matrix lead to the asymptotically valid inference reported by standard statistical software packages. For exact inference, however, the computational burden is heavier: evaluating the validified relative likelihood over a sufficiently fine grid of $\theta$ values is rather expensive. Figure 1(c) presents the $0.1$-level set of the exact IM possibility contour for the regression coefficients, based on a naive Monte Carlo implementation of (2), alongside the proposed variational approximation. The variational method is nearly 2.5 times faster than the naive one, yet the two contours align almost perfectly.
@@
Poisson log-linear models are widely used to analyze the relationship between a count-based discrete response variable and a set of fixed explanatory variables; even if the explanatory variables are not fixed by design, it is almost always assumed that their distribution does not depend on any of the relevant parameters; this makes the explanatory variables ancillary statistics and hence it is customary to condition on their observed values. Let $X_{i}$ represent the response variable, and $z_{i1},\ldots,z_{ip}$ denote the $p$ explanatory variables for the $i^{th}$ observation in a sample, $i=1,\ldots,n$. The Poisson log-linear model assumes a Poisson model for $X_{i}$, given $z_{i}=(z_{i1},\ldots,z_{ip})$, independent across $i=1,\ldots,n$, with marginal mass function
$p_{\lambda_{i}}(x_{i})=\lambda_{i}^{x_{i}}\,e^{-\lambda_{i}}\,/\,x_{i}!,\quad x_{i}=0,1,2,\ldots,$
where $\log\lambda_{i}=\theta_{0}+\theta_{1}z_{i1}+\theta_{2}z_{i2}+\ldots+\theta_{p}z_{ip}.$ The goal is to quantify uncertainty about the unknown true regression coefficients $\Theta_{0},\Theta_{1},\ldots,\Theta_{p}$.
@@
To highlight the IM’a calibration guarantees and provide a comparison with an alternative framework, we conduct a simulation study following the setup described above. We generate 500 datasets, each with $n=25$ observations. The 25 pairs of explanatory variables remain fixed across all datasets and are obtained through a random selection from the 173 pairs in the original dataset described above. Importantly, the selected explanatory variables are scaled so that $\sum_{j=1}^{p}z_{ij}=0$ and $p^{-1}\sum_{j=1}^{p}z_{ij}^{2}=1$—this scaling does not affect the dependence between the explanatory variables, it ensures that the $\Theta_{j}$’s are comparable, so that a hypthesis like $H_{3}$ below is meaningful. The response variable, the $X_{i}$’s, are independently sampled from ${\sf Pois}(\Lambda_{i})$, with
$\log\Lambda_{i}=\Theta_{0}+\Theta_{1}z^{*}_{i1}+\Theta_{2}z^{*}_{i2},$
where $z^{*}_{i1}$ and $z^{*}_{i2}$ represent the scaled $z_{i1}$ and $z_{i2}$, respectively, and $\Theta_{0}=1$, $\Theta_{1}=0.25$ and $\Theta_{2}=0.1$. Consider the following three practically relevant hypotheses,
$H_{1}:\Theta_{1}<0.5;\quad\quad H_{2}:0<\Theta_{2}<0.5;\quad\quad H_{3}:\Theta_{1}>\Theta_{2},$
concerning different features of $\Theta$. Note that all three hypotheses are true, i.e., the true $\Theta$ satisfies the constraints specified by these hypotheses. Figure 3 plots the distribution function of the possibility assigned to these hypotheses by the IM, obtained via the variational approximation described above:
$G_{k}^{\text{\sc im}}(\alpha)=\mathsf{P}_{\Theta}\{\overline{\Pi}_{X^{n}}(H_{k})\leq\alpha\},\quad\alpha\in(0,1),\quad k=1,2,3.$
As expected, the IM possibility assignments are well calibrated for all three hypotheses—(6) implies that this curve would be below the diagonal line. Figure 3 also shows the distribution functions,
$G_{k}^{\text{\sc bayes}}(\alpha)=\mathsf{P}_{\Theta}\{\mathsf{Q}_{X^{n}}(H_{k})\leq\alpha\},\quad\alpha\in(0,1),\quad k=1,2,3,$
of the posterior probabilities $\mathsf{Q}_{X^{n}}(H_{k})$ assigned to the above hypotheses using a Bayesian Poisson regression model, where independent Gaussian priors with mean zero and variance $100^{2}$ were placed on the regression coefficients. While the posterior probabilities are well calibrated for $H_{1}$, it is not calibrated for $H_{2}$ and $H_{3}$—being above the diagonal line means the posterior probabilities assigned to these true hypotheses are stochastically smaller than ${\sf Unif}(0,1)$. In particular, $H_{3}$ gets assigned posterior probabilities less than 0.1 around 20% of the time. This lack of calibration is undesirable, as it can lead to “systematically misleading conclusions” (Reid and Cox, 2015).
Figure 3: Distribution functions $G_{k}^{\text{\sc im}}$ (dashed) and $G_{k}^{\text{\sc bayes}}$ (solid) of the IM possibilities and Bayes posterior probabilities assigned to $H_{1}:\Theta_{1}<0.5$ (blue), $H_{2}:0<\Theta_{2}<0.5$ (green) and $H_{3}:\Theta_{1}>\Theta_{2}$ (red), based on 500 Monte Carlo samples.